OpenStudy (anonymous):

Use the ratio test to determine whether each series is convergent or divergent. a) 2/5 + 4/10 + 8/15 + ...

OpenStudy (anonymous):

\(a_n=\frac{2^n}{5n}\), so the ratio test is: \[\lim_{n\rightarrow \infty}\left|\frac{\frac{2^{n+1}}{5(n+1)}}{\frac{2^n}{5n}}\right|\\ \lim_{n\rightarrow \infty}\left|\frac{2^{n+1}\cdot 5n}{2^n\cdot 5(n+1)}\right|\\ \lim_{n\rightarrow\infty}\frac{2n}{n+1}\\ 2 > 1\] Hence the series diverges.